# Causal Panel Analysis under Parallel Trends:

### v5

@@ Abstract
The finest estimand is the ITE, $\tau_{i,t}=Y_{i,t}(1)-Y_{i,t}(0)$, of which there exists one for each observation $(i,t)$. [^7] This is without loss of generality when feedback and interference are excluded. In staggered DID designs, carryover effects are permissible. When potential outcomes are defined in terms of treatment history, $\tau_{i,t}$ is defined as $Y_{i,t}(1)-Y_{i,t}(\infty)$ where $Y_{i,t}(\infty)$ signifies the untreated potential outcome when unit $i$ never undergoes treatment. Most political science research, however, typically focuses on estimating a single summary statistic. Commonly, this is the ATT, which represents ITE averaged over all observations exposed to the treatment condition. In between these extremes of granularity and coarseness are time-varying dynamic treatment effects (DTE), which are across-unit averages of ITE at each time period relative to treatment adoption (e.g., all observations immediately proceeding treatment adoption). In the staggered adoption setting, we can further subdivide by cohort. We use $\tau_{l}$ ($\tau_{g,l}$) to denote the DTE $l$ periods after treatment adoption (for treatment cohort $g$), with $l=1$ representing the period immediately after treatment adoption. [^8] Some of the authors we reference denote this first post-treatment period with $l=0$. $\tau_{g,l}$ is also what some authors refer to as a cohort average treatment effect on the treated (CATT) (Strezhnev, 2018; Sun and Abraham, 2021) or group-time average treatment effect (Callaway and Sant’Anna, 2021).
Each of the estimators we discuss can be used to estimate $\tau_{l}$. The outcome model analogous to TWFE for estimating DTE is a lags-and-leads specification. For simplicity, we first describe the staggered setting. Let $K_{i,t}=(t-E_{i,t}+1)$ be the number of periods until (when $K_{i,t}\leq 0$) or since unit $i$’s event time at time $t$ (e.g., $K_{i,t}=1$ if unit $i$ switches into treatment at time $t$). Consider a regression based on the following specification:
@@
where $a$ and $b$ are the number of lag and lead terms (BJS, 2024). In the social science literature, the typical practice is to exclude $l=0$, which corresponds to the time period immediately before the transition into the treatment phase, and use it as a reference period (Roth, 2022). Conventionally, $\hat{\tau}^{TWFE}_{l}$ is interpreted as an estimate of $\tau_{l}$ or as a meaningful weighted average of pertinent ITE. Meanwhile, $\hat{\tau}^{TWFE}_{b+}$ is viewed as an estimate for the long-term effect.
HTE-Robust Estimators
In this section, we offer a brief overview and comparison of several recently introduced HTE-robust estimators. We use the term “HTE” to refer to ITE that are nonidentical, i.e. $\tau_{i,t}\neq\tau_{j,s}$ for some $i,j,s,t$. HTE-robust estimators are defined as those that produce causally interpretable estimates under their respective identifying assumptions. For a more comprehensive discussion on these estimators, please refer to the SM.
Summary of HTE-robust estimators.
Table 1 summarizes the estimators we discuss in this paper. The primary difference resides in the mechanics of their estimation strategies: There are methods based on canonical DIDs and methods based on imputation. We refer to the former as DID extensions and the latter as imputation methods. DID extensions use DTE, estimated from local, $2\times 2$ DIDs between treated and control observations, as building blocks. Imputation methods use ITE, estimated as the difference between an imputed outcome under control and the observed outcome (under treatment), as building blocks. The imputation estimator we use in this paper employs TWFE, fitted with all observations under the control condition, to impute treated counterfactuals. Different strategies also entail different assumptions. Each DID extension, for example, relies on a particular type of PT assumption, whereas imputation methods presuppose a TWFE model for untreated potential outcomes and require a zero mean for the error terms.
Another noteworthy difference lies in the settings in which these estimators are applicable: Some estimators function exclusively in settings with staggered treatment adoption, while others can accommodate scenarios with treatment reversals. In the latter setting, all estimators we discuss also require no anticipation and no or limited carryover effects. Furthermore, the estimators diverge in terms of (1) how they select untreated observations as controls for treated units, (2) how they incorporate pre-treatment or exogenous covariates, and (3) the choice of the reference period. We discuss these details further below and in Section A.1 of the SM.
@@
Research article Sun and Abraham (2021) Callaway and Sant’Anna (2021) Baker, Larcker and Wang (2022) de Chaisemartin and D’Haultfœuille (2020) IKW (2023) BJS (2024) LWX (2024)
@@
Comparison group Never-treated or last-treated Never-treated or not-yet-treated Never-treated Matched stable group (not-yet-treated) Matched stable group (not-yet-treated) Imputed Counterfactual (not-yet-treated) Imputed Counterfactual (not-yet-treated)
Reference period(s) Period 0 An arbitrary pre-treatment period Period 0 Untreated period Period 0 All pre-treatment periods All pre-treatment periods
Covariate adjustment Possible extension Outcome & propensity score modeling Outcome modeling Possible extension Refined matched set & outcome modeling Outcome modeling Outcome modeling
DID extensions.
DID extensions are all built from local, $2\times 2$ DID estimates. The overarching strategy for these estimators is to estimate the DTE, $\tau_{l}$ (or $\tau_{g,l}$ for each cohort $g$ in the staggered setting), for each period since the most recent initiation of treatment, $l$, using one or more valid $2\times 2$ DIDs. By “valid,” we mean that the DID includes (1) a pre-period and a post-period and (2) a treated group and a comparison group. The pre-period is such that all observations in both groups are in control, while the post-period is such that observations from the treated group are in treatment and those from the comparison group are in control. The choice of the comparison group is the primary distinction between estimators in this category. To obtain higher-level averages such as the ATT, we then average over our estimates of $\tau_{l}$ (or $\tau_{g,l}$), typically employing appropriate, convex weights.
@@
Stacked DID is another related estimator sometimes used to address HTE concerns. As described by Baker, Larcker and Wang (2022), it involves creating separate sub-datasets for each treated cohort by combining data from that cohort (around treatment adoption) and data from the never-treated cohort from the same periods. These cohort-specific datasets are then “stacked” to form a single dataset. An event study regression akin to Model (2) with the addition of sub-dataset specific unit and time dummies is then run. This method uses the same comparison group as IW and the never-treated version of CSDID without covariates, but stacked DID estimates a single DTE for a given relative period rather than separate estimates for each cohort. Essentially, stacked DID is a special case of IW that uses immutable weights selected by OLS. We denote the corresponding estimand $\tau_{l}^{vw}$ to reflect the fact that it is variance-weighted. These weights will be convex but in general are not proportional to cohort sizes. Thus, while stacked DID avoids the “negative weighting” problem and meets our criteria for HTE-robustness, its estimands are not the same as $\tau_{l}$ or the ATT. This, however, is arguably a minor issue related to external validity, as the estimands can be interpreted as average treatment effects for populations with a different distribution over ITE than the treated observations.
@@
In settings with treatment reversals, separate groups of researchers have converged on the same strategy for choosing a comparison group: matching treated and control observations that belong to units with identical treatment histories. IKW (2023) suggest one such estimator, PanelMatch, which begins by constructing a “matched set” for each observation $(i,t)$ such that unit $i$ transitions into treatment at time $t$. This matched set includes units that both (1) are not under treatment at time $t$ and (2) share the same treatment history as $i$ for a fixed number of periods leading up to the treatment onset. For each treated observation $(i,t)$ and for every post-period $(t+l-1)$ such that unit $i$ is still under treatment, it then estimates a local DID using the same pre-period $s<t$. The treatment “group” comprises solely of $(i,t)$, and the members of the matched set for $(i,t)$ that are still under control during $t+l-1$ serve as the comparison group. To obtain $\tau_{l}$ for a given $l$, it then averages over the corresponding local DID estimates from all treated observations. IKW (2023) propose incorporating covariates by “refining” matched sets and use $l=0$ as the reference period.
@@
Using a similar strategy, de Chaisemartin and D’Haultfœuille (2020) propose a “multiple DID” estimator, DID_multiple. A notable difference is that they include local DIDs for units leaving the treatment and not just those joining the treatment; when there are no treatment reversals or covariates, DID_multiple is a special case of PanelMatch. The original proposal for DID_multiple also only considers the case where we match on a single period and where $l=1$, but since it has been extented. [^10] The most recent implementation, DIDmultiplegtDYN, allows for estimation of dynamic and long-term effects, bringing DID_multiple closer in functionality to PanelMatch. Consequently, the target estimand is not the ATT but rather an average of the contemporaneous effects of “switching” (i.e., the effect of joining or the negative of the effect of leaving at the time of doing so). In the staggered setting, the PanelMatch estimator aligns with the not-yet-treated version of CSDID (without covariate adjustment). We delve into details on the connections between these three estimators in the SM.
@@
All DID extensions are built using local, $2\times 2$ DIDs, and their assumptions reflect this. Specifically, they each rely on a form of the PT assumption—that is, the expected changes in untreated potential outcomes from one period to the next are equal between the treated and the chosen comparison groups. We defer readers to the SM for a fuller account of each method’s assumptions.
The imputation method.
@@
Imputation estimators do not explicitly estimate local DIDs. Instead, they take the difference between the observed outcome and an imputed counterfactual outcome for each treated observation. The connection to the TWFE model is in the functional form assumption used to impute counterfactual outcomes. Specifically, an imputation estimator first fits a parametric model for the potential outcome under control $Y_{i,t}(0)$—in our case, $Y_{i,t}(0)=X_{i,t}{{}^{\prime}}\beta+\alpha_{i}+\xi_{t}+\varepsilon_{i,t}$—using only control observations $\{(i,t):D_{i,t}=0\}$. It is also through this outcome model that one can adjust for time-varying covariates. Then, it imputes $Y_{i,t}(0)$ for all treated observations $\{(i,t):D_{i,t}=1\}$ using the estimated parameters. Finally, it estimates the ITE, $\tau_{i,t}$, for each treated observation $(i,t)$ by calculating the difference between the observation’s observed outcome $Y_{i,t}=Y_{i,t}(1)$ and its imputed counterfactual outcome $\hat{Y}_{i,t}(0)$. Inference for the estimated $\hat{\tau}_{i,t}$ is possible, although uncertainty estimates need to be adjusted to account for the presence of idiosyncratic errors (e.g., Bai and Ng, 2021). BJS (2024) and LWX (2024) each propose estimators in this category. Each paper proposes a more general framework that nests many models, including TWFE. The latter also introduces several specific imputation estimators, one of which uses the TWFE model, and the authors refer to the resulting estimator as the fixed effect counterfactual estimator, or FEct.
@@
Although DID extensions and imputation methods rely on slightly different identification assumptions, these assumptions usually lead to similar observable implications. Researchers commonly use the presence or absence of pretrends to judge how plausible the PT assumption is. In the classic two-group setting, if there are data from multiple pre-treatment periods, researchers can plot the time series of average outcomes of each group and visually inspect whether they indeed trend together. The intuition is that if PT holds and the average outcome trends of the treated and control groups are indeed parallel in pre-treatment periods when $Y(0)$’s are observed for all units, then it is plausible that PT also holds in the post-treatment periods, when $Y(0)$’s are no longer observable for units in the treatment group. Conversely, differential trends in the pre-treatment periods should make us suspicious of PT. In more complex settings or where we wish to control for observed confounders, researchers often use DTE estimates before and after the onset of treatment, $\tau_{l}$, to construct so-called “event-study plots” to judge the presence of pretrends. If PT holds, then pre-treatment DTE estimates should be around zero. We provide a more thorough discussion and an example of the event-study plot in the next section when we introduce our procedure.
Choice of estimators.
@@
In general, we believe the credibility of identifying assumptions is more important than the choice of estimator. After all, in the staggered setting, when assumptions hold, IW, CSDID, PanelMatch, DID_multiple, and the imputation estimator all converge to the same or a similar estimand. However, there are a few reasons to favor imputation estimators. First, imputation estimators can handle complex settings, including those with treatment reversal—which account for over half of the studies in our sample—and can accommodate time-varying covariates, additional fixed effects, and unit- or group-specific time trends commonly seen in social science research. The imputation estimator connects to TWFE through a shared outcome model, and thus any of the aforementioned modifications to the outcome model can be directly mirrored. DID extensions relates to TWFE through their shared connection to DID in the two-unit, two period setting. Classic DID’s inability to naturally accommodate these complexities limits DID extensions on this front. Just like TWFE, the imputation estimator risks misspecification bias, and adding more redundant terms may significantly increase variance. However, we still consider the added flexibility to be an advantage. Second, imputation estimators are the most efficient under homoskedastic errors (BJS, 2024). [^11] In Section A.3 of the SM, we demonstrate that the imputation estimator tends to yield larger $z$-scores, based on data from our sample. In the majority of cases, the imputation estimator has a smaller standard error, and the difference can be especially dramatic for IW and the never-treated version of CSDID, which often discard the vast majority of untreated observations. Moreover, by using the average of all pre-treatment periods as the reference point rather than a single pre-period, as the default in DID extensions, they provide greater power in hypothesis testing for pretrends. The main drawback of imputation estimators is that their current implementations (either FEct or DID_impute) do not allow for automated adjustment of time-invariant covariates, an advantage offered by CSDID and PanelMatch. Adjusting for pre-treatment characteristics can improve credibility of research, as conditional PT may be more plausible than the unconditional one (Sant’Anna and Zhao, 2020).
Data and Procedure

### v6

@@ 2.2 Causal Estimands
The finest estimand is the individual treatment effect, $\tau_{i,t}=Y_{i,t}(1)-Y_{i,t}(0)$, of which there exists one for each observation $(i,t)$. [^8] This is without loss of generality when feedback and interference are excluded. In staggered DID designs, carryover effects are permissible. When potential outcomes are defined in terms of treatment history, $\tau_{i,t}$ is defined as $Y_{i,t}(1)-Y_{i,t}(\infty)$ where $Y_{i,t}(\infty)$ signifies the untreated potential outcome when unit $i$ never undergoes treatment. Most political science research, however, typically focuses on estimating a single summary statistic. Commonly, this is the ATT, which represents individual treatment effects averaged over all observations exposed to the treatment condition. In between these extremes of granularity and coarseness are time-varying dynamic treatment effects, which are across-unit averages of individual treatment effects at each time period relative to treatment adoption. In the staggered adoption setting, we can further subdivide by cohort. We use $\tau_{l}$ ($\tau_{g,l}$) to denote the dynamic treatment effect $l$ periods after treatment adoption (for treatment cohort $g$), with $l=1$ representing the period immediately after treatment adoption. [^9] Some of the authors we reference denote this first post-treatment period with $l=0$. $\tau_{g,l}$ is also what some authors refer to as a cohort average treatment effect on the treated (Strezhnev, 2018; Sun and Abraham, 2021) or group-time average treatment effect (Callaway and Sant’Anna, 2021).
Each of the estimators we discuss can be used to estimate $\tau_{l}$. The outcome model analogous to TWFE for estimating dynamic effects is a lags-and-leads specification. For simplicity, we first describe the staggered setting. Let $K_{i,t}=(t-E_{i,t}+1)$ be the number of periods until (when $K_{i,t}\leq 0$) or since unit $i$’s event time at time $t$ (e.g., $K_{i,t}=1$ if unit $i$ switches into treatment at time $t$). Consider a regression based on the following specification:
@@
where $a$ and $b$ are the number of lag and lead terms (BJS, 2024). In the social science literature, the typical practice is to exclude $l=0$, which corresponds to the time period immediately before the transition into the treatment phase, and use it as a reference period (Roth, 2022). Conventionally, $\hat{\tau}^{TWFE}_{l}$ is interpreted as an estimate of $\tau_{l}$ or as a meaningful weighted average of pertinent individual treatment effects. Meanwhile, $\hat{\tau}^{TWFE}_{b+}$ is viewed as an estimate for the long-term effect.
3.  HTE-Robust Estimators
In this section, we offer a brief overview and comparison of several recently introduced HTE-robust estimators. We use the term “HTE” to refer to individual treatment effects that are arbitrarily heterogeneous, that is, $\tau_{i,t}\neq\tau_{j,s}$ for some $i,j,s,t$. HTE-robust estimators are defined as those that produce causally interpretable estimates under their respective identifying assumptions. For a more comprehensive discussion on these estimators, please refer to the SM.
3.1 Summary of HTE-Robust Estimators
Table 1 summarizes the estimators we discuss in this article. The primary difference resides in the mechanics of their estimation strategies: there are methods based on canonical DIDs and methods based on imputation. We refer to the former as DID extensions and the latter as imputation methods. DID extensions use dynamic treatment effects, estimated from local, $2\times 2$ DIDs between treated and control observations, as building blocks. Imputation methods use individual treatment effects, estimated as the difference between an imputed outcome under control and the observed outcome (under treatment), as building blocks. The imputation estimator we use in this article employs TWFE, fitted with all observations under the control condition, to impute treated counterfactuals. Different strategies also entail different assumptions. Each DID extension, for example, relies on a particular type of PT assumption, whereas imputation methods presuppose a TWFE model for untreated potential outcomes and require a zero mean for the error terms, which is implied by strict exogeneity.
Another noteworthy difference lies in the settings in which these estimators are applicable: Some estimators can only be used in settings with staggered treatment adoption, while others can accommodate scenarios with treatment reversals. In the latter setting, all estimators we discuss also require no anticipation and no or limited carryover effects. Furthermore, the estimators diverge in terms of (1) how they select untreated observations as controls for treated units, (2) how they incorporate pre-treatment or exogenous covariates, and (3) the choice of the reference period. We discuss these details further below and in Section A.1 of the SM.
@@ 3.1 Summary of HTE-Robust Estimators
Research article Sun and Abraham (2021) Callaway and Sant’Anna (2021) Baker, Larcker and Wang (2022) de Chaisemartin and D’Haultfœuille (2020, 2024) IKW (2023) BJS (2024) LWX (2024)
@@
Comparison group Never-treated or last-treated Never-treated or not-yet-treated Never-treated Matched stable group (not-yet-treated) Matched stable group (not-yet-treated) Imputed counterfactuals (not-yet-treated) Imputed counterfactuals (not-yet-treated)
Reference period(s) Period 0 An arbitrary pre-treatment period Period 0 Period 0 Period 0 All pre-treatment periods All pre-treatment periods
Covariate adjustment Possible extension Outcome & propensity score modeling Outcome modeling Possible extension Refined matched set and outcome modeling Outcome modeling Outcome modeling
3.2 DID Extensions
DID extensions are all built from local, $2\times 2$ DID estimates. The overarching strategy for these estimators is to estimate the dynamic treatment effects, $\tau_{l}$ (or $\tau_{g,l}$ for each cohort $g$ in the staggered setting), for each period since the most recent initiation of treatment, $l$, using one or more valid $2\times 2$ DIDs. By “valid,” we mean that the DID includes (1) a pre-period and a post-period and (2) a treated group and a comparison group. The pre-period is such that all observations in both groups are in control, whereas the post-period is such that observations from the treated group are in treatment and those from the comparison group are in control. The choice of the comparison group is the primary distinction between estimators in this category. To obtain higher-level averages such as the ATT, we then average over our estimates of $\tau_{l}$ (or $\tau_{g,l}$), typically employing appropriate, convex weights.
@@ 3.2 DID Extensions
Stacked DID, first formally introduced by Cengiz et al. (2019), is another related estimator sometimes used to address HTE concerns. As described by Baker, Larcker and Wang (2022), it involves creating separate sub-datasets for each treated cohort by combining data from that cohort (around treatment adoption) and data from the never-treated cohort from the same periods. These cohort-specific datasets are then “stacked” to form a single dataset. An event-study regression akin to Equation (2) with the addition of sub-dataset specific unit and time dummies is then run. This method uses the same comparison group as IW and the never-treated version of CSDID without covariates, but stacked DID estimates a single dynamic treatment effect for a given relative period rather than separate estimates for each cohort. Essentially, stacked DID is a special case of IW that uses immutable weights selected by OLS. We denote the corresponding estimand $\tau_{l}^{vw}$ to reflect the fact that it is variance-weighted. These weights are generally neither proportional to cohort sizes nor guaranteed to sum to one (Wing, Freedman and Hollingsworth, 2024). Thus, while stacked DID avoids the “negative weighting” problem and meets our criteria for HTE-robustness, its estimands are not the same as $\tau_{l}$ or the ATT.
@@
In settings with treatment reversals, separate groups of researchers have converged on the same strategy for choosing a comparison group: matching treated and control observations that belong to units with identical treatment histories. IKW (2023) suggest one such estimator, PanelMatch, which begins by constructing a “matched set” for each observation $(i,t)$ such that unit $i$ transitions into treatment at time $t$. This matched set includes units that both (1) are not under treatment at time $t$ and (2) share the same treatment history as $i$ for a fixed number of periods leading up to the treatment onset. For each treated observation $(i,t)$ and for every post-period $(t+l-1)$ such that unit $i$ is still under treatment, it then estimates a local DID using the same pre-period $s<t$. The treatment “group” comprises solely of observation $(i,t)$, and the members of the matched set for $(i,t)$ that are still under control during $t+l-1$ serve as the comparison group. To obtain $\tau_{l}$ for a given $l$, it then averages over the corresponding local DID estimates from all treated observations. IKW (2023) propose incorporating covariates by “refining” matched sets and use $l=0$ as the reference period.
@@
Using a similar strategy, de Chaisemartin and D’Haultfœuille (2020) propose a “multiple DID” estimator, DID_multiple. A notable difference is that they include local DIDs for units leaving the treatment and not just those joining the treatment; when there are no treatment reversals or covariates, DID_multiple is a special case of PanelMatch. The original proposal for DID_multiple also only considers the case where we match on a single period and where $l=1$, but since it has been extended (de Chaisemartin and D’Haultfœuille, 2024). Consequently, the target estimand is not the ATT but rather an average of the contemporaneous effects of “switching” (i.e., the effect of joining or the negative of the effect of leaving at the time of doing so). In the staggered setting, the PanelMatch estimator aligns with the not-yet-treated version of CSDID (without covariate adjustment). We delve into details on the connections between these three estimators in the SM.
@@
All DID extensions are built using local, $2\times 2$ DIDs, and their assumptions reflect this. Specifically, they each rely on a form of the PT assumption—that is, the expected changes in untreated potential outcomes from one period to the other are equal between the treated and the chosen comparison groups. We defer readers to the SM for a fuller account of each method’s assumptions.
3.3 The Imputation Method
@@
Imputation estimators do not explicitly estimate local DIDs. Instead, they take the difference between the observed outcome and an imputed counterfactual outcome for each treated observation. The connection to the TWFE model is in the functional form assumption used to impute counterfactual outcomes. Specifically, an imputation estimator first fits a parametric model for the potential outcome under control $Y_{i,t}(0)$—in our case, $Y_{i,t}(0)=X_{i,t}{{}^{\prime}}\beta+\alpha_{i}+\xi_{t}+\varepsilon_{i,t}$—using only control observations $\{(i,t):D_{i,t}=0\}$. It is also through this outcome model that one can adjust for time-varying covariates. Then, it imputes $Y_{i,t}(0)$ for all treated observations $\{(i,t):D_{i,t}=1\}$ using the estimated parameters. Finally, it estimates the individual treatment effect, $\tau_{i,t}$, for each treated observation $(i,t)$ by calculating the difference between the observation’s observed outcome $Y_{i,t}=Y_{i,t}(1)$ and its imputed counterfactual outcome $\hat{Y}_{i,t}(0)$. Inference for the estimated $\hat{\tau}_{i,t}$ is possible, although uncertainty estimates need to be adjusted to account for the presence of idiosyncratic errors (e.g., Bai and Ng, 2021). BJS (2024) and LWX (2024) each propose estimators in this category. Each paper proposes a more general framework that nests many models, including TWFE. The latter also introduces several specific imputation estimators, one of which uses the TWFE model, and the authors refer to the resulting estimator as the fixed effects counterfactual estimator, or FEct.
@@
Although DID extensions and imputation methods rely on slightly different identification assumptions, these assumptions usually lead to similar observable implications. Researchers commonly use the presence or absence of pretrends to judge how plausible the PT assumption is. In the classic two-group setting, if there are data from multiple pre-treatment periods, researchers can plot the time series of average outcomes of each group and visually inspect whether they indeed trend together. The intuition is that if PT holds and the average outcome trends of the treated and control groups are indeed parallel in pre-treatment periods when $Y(0)$’s are observed for all units, then it is plausible that PT also holds in the post-treatment periods, when $Y(0)$’s are no longer observable for units in the treatment group. Conversely, differential trends in the pre-treatment periods should make us suspicious of PT. In more complex settings or where we wish to control for observed confounders, researchers often use dynamic estimates before and after the onset of treatment, $\tau_{l}$, to construct so-called “event-study plots” to judge the presence of pretrends. If PT holds, then pre-treatment dynamic estimates should be around zero. We provide a more thorough discussion and an example of the event-study plot in the next section when we introduce our procedure.
3.4 Choice of Estimators
@@
In general, we believe the credibility of identifying assumptions is more important than the choice of estimator. After all, in the staggered setting, when assumptions hold, IW, CSDID, DID_multiple, PanelMatch, and the imputation estimator all converge to the same or a similar estimand. However, there are a few reasons to favor the imputation estimator. First, it can handle complex settings, including those with treatment reversal—which account for over half of the studies in our sample—and can accommodate time-varying covariates, additional fixed effects, and unit- or group-specific time trends commonly seen in social science research. The imputation estimator connects to TWFE through a shared outcome model, and thus any of the aforementioned modifications to the outcome model can be directly mirrored. DID extensions relates to TWFE through their shared connection to DID in the two-group, two period setting. Classic DID’s inability to naturally accommodate these complexities limits DID extensions on this front. Just like TWFE, the imputation estimator risks misspecification bias, and adding more redundant terms may significantly increase variance. However, we still consider the added flexibility to be an advantage. Second, imputation estimators are the most efficient under homoskedastic errors (BJS, 2024). [^11] In Section A.3 of the SM, we demonstrate that the imputation estimator tends to yield larger $z$-scores, based on data from our sample. In the majority of cases, the imputation estimator has a smaller standard error, and the difference can be especially dramatic for IW and the never-treated version of CSDID, which often discard the vast majority of untreated observations. Moreover, by using the average of all pre-treatment periods as the reference point rather than a single pre-period, as the default in DID extensions, they provide greater power in hypothesis testing for pretrends. The main drawback of imputation estimators is that their current implementations (either FEct or DID_impute) do not allow for automated adjustment of time-invariant covariates, an advantage offered by CSDID and PanelMatch. Adjusting for pre-treatment characteristics can improve credibility of research, as conditional PT may be more plausible than the unconditional one (Sant’Anna and Zhao, 2020).
4.  Data and Procedure